Liquid N2O3 is dark blue at low temperatures, but the color fades and becomes greenish at higher temperatures as the compound decomposes to NO and NO2. At 25 degree C, a value of Kp = 1.91 has been established for this decomposition. If 0236 moles of N2O3 are placed in a 1.52-L vessel at 25 degree C, calculate the equilibrium partial pressures of N2O3(g), NO2(g) and NO(g).
Answer #1

N2O3(g) ----> NO2(g) + NO(g)

Kp = pNO*pNO2 / pN2O3

INITIAL PRESSURE OF n2o3 = nRT/V = 0.236*0.0821*298/1.52

   = 3.8 atm

at equilibrium

pN2O3 = 3.8-x

pNO = X

pNO2 = X


1.91 = X^2/(3.8-X)

X = 1.9

partial pressure of N2O3 = 3.8-1.9 = 1.9 atm

partial pressure of NO = 1.9 atm

partial pressure of NO2 = 1.9 atm